# Block-Recurrent Dynamics in ViTs

### v2

@@ 1 Introduction
One promising path toward such understanding is to search for underlying simplicity. Multiple approaches explore different facets of this simplicity, whether in functional expressivity (Montúfar et al., 2014; Telgarsky, 2015; Serra et al., 2018; Balestriero and others, 2018), symmetry (Cohen and Welling, 2014; Olah et al., 2020), or computation (Wilson, 2025; Goldblum et al., 2023; Schmidhuber, 1997; Mingard et al., 2025). Discovering such simplicity principle should improve both development (Bronstein et al., 2017; Frankle and Carbin, 2019) and interpretability (Bereska and Gavves, 2024; Carvalho et al., 2019; Fel, 2024; Ghorbani et al., 2017; Fel et al., 2023; Smilkov et al., 2017; Sundararajan et al., 2017; Zeiler and Fergus, 2014; Templeton et al., 2024; Bricken et al., 2023). Depth offers a concrete place to look for this simplicity. Residual connections have long suggested a link to dynamical systems (Liao and Poggio, 2020; Veit et al., 2016; Greff et al., 2016; Boulch, 2017; Haber and Ruthotto, 2017), hinting at implicit recurrence even when layers have distinct parameters. This convergent evidence makes plausible a form of algorithmic parsimony (Ma et al., 2022) in which a small set of blocks may be reused across many layers, trading parameters for iterations. Related perspectives support this view (Dingle et al., 2020). More concretely, residual updates invite a discrete-time interpretation of depth (Chen et al., 2018; Chalvidal et al., 2020; Sander et al., 2022), attention induces coupled token dynamics (Lu et al., 2019; Geshkovski et al., 2023), and in language models contiguous block recurrence has been observed and exploited (Geiping et al., 2025; Fernando and Guitchounts, 2025; Dehghani et al., 2018; Tan et al., 2023). This renewed interest in recurrent computation (Jolicoeur-Martineau, 2025; Venkatraman et al., 2025) and algorithmic complexity (Shaw et al., 2025; Dingle et al., 2020) as frameworks for understanding neural network simplicity bias further motivates our investigation.
@@ Results.
Further gains come from depth scaling, where each block has a learned vector embedding of its target layer index, making Raptor a non-autonomous dynamical system (i.e., the update rule explicitly depends on the iteration count rather than state alone). Additional improvements come from up-weighting the cls token loss in the final block (see Eq. A, Appendix A). Finally, connecting all blocks and fine-tuning the model end-to-end with the autoregressive objective (i.e. the second stage) produces a dramatic jump in performance, and a final boost is obtained by fine-tuning the linear probe. Now that we have shown that the BRH holds for a foundation model, and before turning to Dynamical Interpretability, we first examine one implication of this phenomena: the algorithmic and computational implications of the BRH.
@@ Appendix A Training Block Recurrent Foundation Models
All Raptor variants ($k=2,3,4$) are trained on top of DINOv2 (ViT-B with registers) (Darcet et al., 2023) and use the same transformer architecture. The model utilizes a feature dimension of $784$, an MLP ratio of $4$, and $12$ multi-head attention heads. The depth-scale MLP consists of a linear layer expanding from dimension 1 to 16, followed by a SiLU activation, and a second linear layer mapping from 16 to $3\times\text{dim}$. This produces three separate scaling vectors. For each Raptor  variant in Table 1 and Figure 7, we train train three models, each with a different seed. We measure the performance of each model and report aggregate statistics. Below, we provide the hyperparameters and training settings used to train Raptor, describing the layer divisions for different values of $k$ followed by the training procedure.
@@ A.2 Teacher-student reconstruction $R^{2}$
A.3 Linear Probe Fine-tuning

### v4

@@ 1 Introduction
One promising path toward such understanding is to search for underlying simplicity. Multiple approaches explore different facets of this simplicity, whether in functional expressivity (Montúfar et al., 2014; Telgarsky, 2015; Serra et al., 2018; Balestriero and others, 2018), symmetry (Cohen and Welling, 2014; Olah et al., 2020), or computation (Wilson, 2025; Goldblum et al., 2023; Schmidhuber, 1997; Mingard et al., 2025). Discovering such simplicity principle should improve both development (Bronstein et al., 2017; Frankle and Carbin, 2019) and interpretability (Bereska and Gavves, 2024; Carvalho et al., 2019; Fel, 2024; Ghorbani et al., 2017; Fel et al., 2023; Smilkov et al., 2017; Sundararajan et al., 2017; Zeiler and Fergus, 2014; Templeton et al., 2024; Bricken et al., 2023). Depth offers a concrete place to look for this simplicity. Residual connections have long suggested a link to dynamical systems (Liao and Poggio, 2020; Veit et al., 2016; Greff et al., 2016; Boulch, 2017; Haber and Ruthotto, 2017), hinting at implicit recurrence even when layers have distinct parameters. This convergent evidence makes plausible a form of algorithmic parsimony (Ma et al., 2022) in which a small set of blocks may be reused across many layers, trading parameters for iterations (Schwarzschild et al., 2021). Related perspectives support this view (Dingle et al., 2020). More concretely, residual updates invite a discrete-time interpretation of depth (Chen et al., 2018; Chalvidal et al., 2020; Sander et al., 2022), attention induces coupled token dynamics (Lu et al., 2019; Geshkovski et al., 2023), and in language models contiguous block recurrence has been observed and exploited (Geiping et al., 2025; Fernando and Guitchounts, 2025; Dehghani et al., 2018; Tan et al., 2023). This renewed interest in recurrent computation (Jolicoeur-Martineau, 2025; Venkatraman et al., 2025) and algorithmic complexity (Shaw et al., 2025; Dingle et al., 2020) as frameworks for understanding neural network simplicity bias further motivates our investigation.
@@ Results.
Further gains come from depth scaling (Appendix A.3), where each block has a learned vector embedding of its target layer index, making Raptor a non-autonomous dynamical system (i.e., the update rule explicitly depends on the iteration count rather than state alone). Additional improvements come from up-weighting the cls token loss in the final block (see Eq. A, Appendix A). Finally, connecting all blocks and fine-tuning the model end-to-end with the autoregressive objective (i.e. the second stage) produces a dramatic jump in performance, and a final boost is obtained by fine-tuning the linear probe. Now that we have shown that the BRH holds for a foundation model, and before turning to Dynamical Interpretability, we first examine one implication of this phenomena: the algorithmic and computational implications of the BRH.
@@ Appendix A Training Block Recurrent Foundation Models
All Raptor variants ($k=2,3,4$) are trained on top of DINOv2 (ViT-B with registers) (Darcet et al., 2023) and use the same transformer architecture. The model utilizes a feature dimension of $784$, an MLP ratio of $4$, and $12$ multi-head attention heads. The depth-scale MLP in each block generates three separate scaling vectors; we provide the full formulation in Appendix A.3. For each Raptor  variant in Table 1 and Figure 7, we train train three models, each with a different seed. We measure the performance of each model and report aggregate statistics. Below, we provide the hyperparameters and training settings used to train Raptor, describing the layer divisions for different values of $k$ followed by the training procedure.
@@ A.2 Teacher-student reconstruction $R^{2}$
A.3 Depth-Scaling Mechanism
We employ a lightweight conditioning MLP in each block that maps a scalar depth coordinate $z\in\mathbb{R}$ to layer-specific scaling vectors. For a layer with feature dimension $D=784$, the depth-scale MLP generates a modulation vector $\mathbf{v}\in\mathbb{R}^{3D}$.
The MLP projects the scalar input to a hidden dimension $h=16$, applies a SiLU activation, and projects to the output dimension. We add a unity offset to the MLP output to ensure the scaling factors center around 1 (identity behavior) at initialization. The compound scaling vector $\mathbf{S}$ is computed as:
$\mathbf{S}=\left(\mathbf{W}_{2}\cdot\text{SiLU}(\mathbf{W}_{1}z+\mathbf{b}_{1})+\mathbf{b}_{2}\right)+\mathbf{1}$ (4)
where $\mathbf{1}$ denotes a vector of ones. We initialize the weights and biases of the final projection ($\mathbf{W}_{2},\mathbf{b}_{2}$) with a normal distribution ($\mu=0,\sigma=10^{-4}$), ensuring that the block initially acts close to standard behavior.
The vector $\mathbf{S}$ is split along the feature dimension into three component vectors: $\mathbf{s}_{\text{attn}},\mathbf{s}_{\text{mlp}},\mathbf{s}_{\text{out}}\in\mathbb{R}^{D}$. These vectors modulate the signal via element-wise multiplication ($\odot$) at three points in the transformer block: the attention residual, the MLP residual, and the final block output. Given an input sequence $\mathbf{X}$, the forward pass is defined as:
$\mathbf{X}^{\prime}$ $=\mathbf{X}+\mathbf{s}_{\text{attn}}\odot\text{LS}_{1}(\text{Attn}(\text{LN}_{1}(\mathbf{X})))$ (5)
$\mathbf{X}^{\prime\prime}$ $=\mathbf{X}^{\prime}+\mathbf{s}_{\text{mlp}}\odot\text{LS}_{2}(\text{MLP}(\text{LN}_{2}(\mathbf{X}^{\prime})))$ (6)
$\mathbf{Y}$ $=\mathbf{s}_{\text{out}}\odot\mathbf{X}^{\prime\prime}$ (7)
where LS refers to LayerScale, LN is Layer Normalization, and broadcasting is applied over the sequence length dimension.
A.4 Linear Probe Fine-tuning